OpenStudy (anonymous):

What is the square root of 500x^3 all over the square root of 10x^-1 ??

OpenStudy (lgbasallote):

\[\huge \frac{\sqrt{500 x^3}}{\sqrt{10x^{-1}}} \implies \sqrt{\frac{500 x^3}{10x^{-1}}}\implies \sqrt{\frac{500x^3 \cdot x}{10}}\] does that help?

OpenStudy (anonymous):

do i need to simplify this ??

OpenStudy (lgbasallote):

of course.

OpenStudy (anonymous):

wow that is large

OpenStudy (anonymous):

okay so i just divide and then what do i do with the exponents my teacher hasnt reviewed tthis yet but it was on homework